January 1997

Sunday  Monday  Tuesday  Wednesday  Thursday  Friday  Saturday

1

New Year

10:00 Wednesday call 16:00 Nancy 18:00–20:00 Dinner

2 3

Third of month

10:00–13:00 Koffee klatch

4

5 6

Monday event 14:00–16:00 Monday meeting

7 8 9 10

Alternate Monday

11

12 13

Monday event 14:00–16:00 Monday meeting

14 15

10:00 Wednesday call

16 17

18

19 20

Monday event 14:00–16:00 Monday meeting

21 22 23 24

Alternate Monday

25

26 27

Monday event 14:00–16:00 Monday meeting

28 29

10:00 Wednesday call

30 31